OpenStudy (skittles_for_life6422):

What are the minimum, first quartile, median, third quartile, and maximum of the data set? 18, 20, 11, 10, 8, 6, 12, 4

OpenStudy (skittles_for_life6422):

answer choices: minimum 4; first quartile 7; median 10.5; third quartile 17.5; maximum 20 minimum 4; first quartile 5.5; median 12.75; third quartile 15; maximum 20 minimum 4; first quartile 8.75; median 12.75; third quartile 17.5; maximum 20 minimum 4; first quartile 7; median 10.5; third quartile 15; maximum 20

OpenStudy (skittles_for_life6422):

@Hero @dan815 @nincompoop @Michele_Laino

OpenStudy (skittles_for_life6422):

i think its a or d

OpenStudy (anonymous):

SKITTLES

OpenStudy (anonymous):

its a

OpenStudy (skittles_for_life6422):

are you sure can you show why

OpenStudy (anonymous):

first we put the numbers in order... 4,6,8,10,11,12,18,20 now the minimum is the lowest number and the maximum is the highest number. the median is the middle number...if there is 2 middle numbers, add them then divide by 2 ..so median is : (10 + 11) / 2 = 21/2 = 10.5 Q1 is the middle number of all the numbers below the median 4,6,8,10....middle number is (6 + 8)/2 = 7 Q3 is the middle number of all the numbers above the median 11,12,18,20.....(12 + 18)/2 = 15 so answer is D

OpenStudy (skittles_for_life6422):

that makes sense thanks for explaining

OpenStudy (anonymous):

no problem :)

OpenStudy (skittles_for_life6422):

can you help with 6 more questions

OpenStudy (anonymous):

possibly...as long as they are not geometry...not too good at that

OpenStudy (skittles_for_life6422):

i'll just post them and we'll see if you can explain them

OpenStudy (anonymous):

sounds good

OpenStudy (skittles_for_life6422):

simplify. \[-(2xy ^{3)^{-5}}\]

OpenStudy (skittles_for_life6422):

answer choices: \[-32x ^{5}y ^{15}\] \[\frac{ -32 }{ x ^{5}y ^{15}}\] \[\frac{ -1 }{ 32x ^{5}y ^{15} }\] \[\frac{ -1 }{ 32x ^{4}y ^{2} }\]

OpenStudy (anonymous):

-(2xy^3)^-5 = - 1/(2xy^3)^5 = - 1/(32x^5y^15)

OpenStudy (skittles_for_life6422):

thanks here is the next one

OpenStudy (skittles_for_life6422):

what is the simplified form of the expression \[4d ^{-3} \times d^{18} \]

OpenStudy (skittles_for_life6422):

answer choices: \[5d ^{-54}\] \[5d ^{15}\] \[4d ^{15}\] \[4d ^{-54}\]

OpenStudy (anonymous):

4d^-3 * d^18 = 4d^15

OpenStudy (skittles_for_life6422):

ok here is the next one

OpenStudy (skittles_for_life6422):

What is the degree of the monomial. \[6x ^{5}\] answer choices: 6 30 5 11

OpenStudy (anonymous):

the degree of the monomial is the sum of the exponents of all the variables...and since there is only 1 variable, and its exponent is 5...then it is a degree of 5

OpenStudy (skittles_for_life6422):

solve the equation by completing the square. Round to the nearest hundredth if necessary. \[x ^{2}-6x=7\] answer choices: x=7,1 x=-7,1 x=-7,-1 x=7,-1

OpenStudy (anonymous):

SHUT UP I am trying to ask questions and i am getting millions of notifications from this

OpenStudy (anonymous):

MESSAGE EACHOTHER ALONE

OpenStudy (anonymous):

x^2 - 6x = 7 take half of coefficient and square it -3^2 = 9 add that to both sides x^2 - 6x + 9 = 7 + 9 (x - 3)^2 = 16 x - 3 = (+-) sq rt 16 x = 3 (+-) 4 x = 3 + 4 x = 7 x = 3 - 4 x = -1 answer is 7 and -1

OpenStudy (anonymous):

I am sorry but I have to go now...I wish you the best of luck

OpenStudy (skittles_for_life6422):

but i only have 2 more

OpenStudy (anonymous):

I am sorry....I am already gonna be late...I lost track of time....soooo sorry

OpenStudy (anonymous):

SHUT UP

OpenStudy (anonymous):

sorry cooljosh :(

OpenStudy (anonymous):

ITS OKAY

OpenStudy (anonymous):

oops caps lock was still on :(

OpenStudy (skittles_for_life6422):

solve the equation. identify any extraneous solutions answer choice: 16 is a solution to the original equation 16 is a solution to the original equation.-16 is an extraneous solution. no solution -16 is a solution to the original equation. 16 is an extraneous solution